# Quantum many-body simulations with PauliStrings.jl

### v2

@@ II Methods
$\tau_{ij}\tau_{kl}=\alpha\tau_{(i\oplus k)(j\oplus l)}.$ (6)
We can use this property to efficiently multiply Pauli strings. Encode a Pauli string $\tau_{a}$ in a tuple of binary integers $a=(v,w)$ such that $\tau_{a}=\tau_{v_{1}w_{1}}\otimes\tau_{v_{2}w_{2}}\otimes\tau_{v_{3}w_{3}}...$ where $v_{i}$ is the $i^{th}$ bit of integer $v$. Then the following relation holds:
$(-1)^{\alpha_{1}}\tau_{a_{1}}(-1)^{\alpha_{2}}\tau_{a_{2}}=(-1)^{\alpha_{12}}\tau_{a_{12}},$ (7)
@@
$\alpha_{12}$ $=(-1)^{\textup{pop}(v_{1}\wedge w_{2})}.$ (10)
@@
$\left[(-1)^{\alpha_{1}}\tau_{a_{1}},(-1)^{\alpha_{2}}\tau_{a_{2}}\right]=(-1)^{\alpha_{12}}\tau_{a_{12}}$ (11)
@@
$\alpha_{12}=(-1)^{\textup{pop}(v_{1}\wedge w_{2})}-(-1)^{\textup{pop}(w_{1}\wedge v_{2})}.$ (12)
@@
Algorithm 1 Product of two operators $A$ and $B$ in the binary Pauli string representation
$C\leftarrow$ empty operator
for $0\leq i<$length($A$do
     for $0\leq j<$length($B$do
         $v\leftarrow A.v_{i}\oplus B.v_{j}$
         $v\leftarrow A.w_{i}\oplus B.w_{j}$
         $h\leftarrow A.h_{i}\cdot B.h_{j}\cdot(-1)^{\textup{pop}(A.v_{i}\wedge B.w_{j})}$
@@
In the special case that the operators have translation symmetry, there is an even more efficient way to encode them. Consider for example the 1D Ising Hamiltonian with periodic boundary conditions $H=-J(\sum_{i}Z_{i}Z_{i+1}+g\sum_{i}X_{i})$. In this case, there is no need to store all the Pauli strings. $H$ is fully specified by the two strings $-JZ_{1}Z_{2}$ and $-JgX_{1}$ and the fact that it has translation symmetry. In general, a 1D translation symmetric operator can be written as $\sum_{i}T_{i}(H_{0})$ where $T_{i}$ is the $i$-site’s translation operator and $H_{0}$ is the local operator that generates $H$. $H_{0}$ can be chosen so that it’s only composed of Pauli strings that start on the first site. Algorithm 2 shows how to take the product of two translation symmetric operators. The main difference from Algorithm 1 is that we need to translate each string back so that it starts on the first site. In PauliStrings.jl, this is implemented in the OperatorTS1D structure.
@@
Algorithm 2 Product of two translation symmetric operators $A$ and $B$. $T_{k}$ is the translation operator by $k$ sites and Shiftleft($\tau$) translates $\tau$ such that it starts on the first site.
@@
[…]
              $h\leftarrow A.h_{i}\cdot B.h_{j}$
              if $\tau$ is in $C$ then
                  $C$[$\tau$] $\leftarrow$ $C$[$\tau$]+$h$
@@
Matrix Product States and Matrix Product Operators are 1D tensor networks in which $N$ spins are represented by $N$, 3-dimensional tensors, each with two internal indices running over the internal ‘bond-dimension’ (BD), and the third index representing the spin degrees of freedom. When doing calculations a maximum bond-dimension of tensors can be set, inducing a controlled error in the representation of the state/operator, but at the benefit of reducing the number of parameters to be just polynomial in the system size.
@@ III Heisenberg time evolution
using the Runge-Kutta method. To keep the number of strings manageable, we introduce noise and truncate the operator $O$ at each time step by keeping the strings with the largest weight. The noise is modeled by a depolarizing channel that causes long Pauli strings to decay. In the Heisenberg picture, observables evolve under the adjoint channel. Because a depolarizing channel is self-adjoint, we can apply it directly to $O$. The transmission rate associated with a Pauli string of length $w$ is $e^{-\epsilon w}$ where $\epsilon$ is the noise amplitude. Similar approaches have been recently used in Refs.  [11, 20, 21, 22, 23]. The strategy here is to choose the smallest noise value that makes the simulations tractable, while not destroying the phenomena of interest.
@@ IV Krylov subspace expansion
Lets introduce Lanczos’ algorithm. The idea is to construct an orthonormal basis of operators generated by recursively applying $\mathcal{L}$ to $O$. First, one defines an inner product, which here we choose to be the infinite temperature inner product: $AB=\frac{1}{2^{N}}{\rm Tr}[A^{\dagger}B]$ and norm $\lVert O\rVert=\frac{1}{2^{N}}{\rm Tr}[O^{2}]$. Then, the first iteration is given by:
@@ Acknowledgements.
N.L. was supported by a research grant (42085) from Villum Fonden. N.L. thanks Prof. Berislav Buca for funding and support, which made finishing this project possible. D.S. and J.C.P. are partially supported by AFOSR (Award no. FA9550-21-1-0236) and NSF (Award no. OAC-2118310). This work was supported in part through the NYU IT High Performance Computing resources, services, and staff expertise.

### v3

@@ II Methods
$\tau_{v_{1}w_{1}}\tau_{v_{2}w_{2}}=\alpha\tau_{(v_{1}\oplus v_{2})(w_{1}\oplus w_{2})}.$ (6)
We can use this property to efficiently multiply Pauli strings. Encode a Pauli string $\tau_{a}$ in a tuple of binary integers $a=(v,w)$ such that $\tau_{a}=\tau_{v_{(1)}w_{(1)}}\otimes\tau_{v_{(2)}w_{(2)}}\otimes\tau_{v_{(3)}w_{(3)}}...$ where $v_{(i)}$ is the $i^{th}$ bit of integer $v$. Then the following relation holds:
$\tau_{a_{1}}\tau_{a_{2}}={\alpha}\tau_{a_{12}},$ (7)
@@
$\alpha$ $=(-1)^{\textup{pop}(v_{1}\wedge w_{2})}.$ (10)
@@
$\left[\tau_{a_{1}},\tau_{a_{2}}\right]=\alpha\tau_{a_{12}}$ (11)
@@
$\alpha=(-1)^{\textup{pop}(v_{1}\wedge w_{2})}-(-1)^{\textup{pop}(w_{1}\wedge v_{2})}.$ (12)
@@
Algorithm 1 Product of two operators $A$ and $B$ in the binary Pauli string representation. Operators are encoded as dictionaries of complex numbers indexed by tuple of integers $(v,w)$ that represent a Pauli string. In this algorithm, $A_{v,w}$ denotes the coefficient in front of the string encoded by $v,w$ in operator $A$.
$C\leftarrow$ empty dictionary
for $(v_{A},w_{A})$ in $A$.keys do
     for $(v_{B},w_{B})$ in $B$.keys do
         $v\leftarrow v_{A}\oplus v_{B}$
         $w\leftarrow w_{A}\oplus w_{B}$
         $h\leftarrow A_{v_{A},w_{A}}\cdot B_{v_{B},w_{B}}\cdot(-1)^{\textup{pop}(v_{A}\wedge w_{B})}$
@@
In the special case that the operators have translation symmetry, there is an even more efficient way to encode them. Consider for example the 1D Ising Hamiltonian with periodic boundary conditions $H=-J(\sum_{i}Z_{i}Z_{i+1}+g\sum_{i}X_{i})$. In this case, there is no need to store all the Pauli strings. $H$ is fully specified by the two strings $-JZ_{1}Z_{2}$ and $-JgX_{1}$ and the fact that it has translation symmetry. In general, a 1D translation symmetric operator can be written as $\sum_{i}T_{i}(H_{0})$ where $T_{i}$ is the $i$-site’s translation operator and $H_{0}$ is the local operator that generates $H$. $H_{0}$ can be chosen so that it’s only composed of Pauli strings that start on the first site. Algorithm 2 shows how to take the product of two 1D translation symmetric operators. The main difference from Algorithm 1 is that we need to translate each string back so that it starts on the first site. In PauliStrings.jl, this is implemented in the OperatorTS1D structure. Note that Algorithm 2 can easily be adapted to higher dimensions by iterating over the necessary shifts corresponding to each dimension.
@@
Algorithm 2 Product of two translation symmetric operators $A$ and $B$ supported on $N$ spins. $T_{k}$ is the translation operator by $k$ sites and Shiftleft($v,w$) translates the string $(v,w)$ such that it starts on the first site.
@@
              $(v_{B}^{\prime},w_{B}^{\prime})\leftarrow T_{k}(v_{B},w_{B})$
              $(v,w)\leftarrow(v_{A}\oplus v_{B}^{\prime},w_{A}\oplus w_{B}^{\prime})$
              $(v,w)\leftarrow$Shiftleft$(v,w)$
              $h\leftarrow A_{v_{A},w_{A}}\cdot B_{v_{B},w_{B}}\cdot(-1)^{\textup{pop}(v_{A}\wedge w_{B})}$
              if $(v,w)$ is in $C$ then
                  $C_{v,w}$leftarrow $C_{v,w}$+$h$
@@
Matrix Product States (MPS) are 1D tensor networks which can be used to represent the quantum state of $N$ spins by constructing $N$, 3-dimensional tensors; two of the indices run over an internal ‘bond-dimension’ (BD) and the third index represents the spin degrees of freedom. Matrix Product Operators (MPO) are similarly used to represent quantum operators and consist of $N$ tensors, each with 4-dimensions (two bond-dimensions and two spin degrees of freedom). When doing calculations a maximum bond-dimension of tensors can be set, inducing a controlled error in the representation of the state/operator, but at the benefit of reducing the number of parameters to be just polynomial in the system size.
@@ III Heisenberg time evolution
using the Runge-Kutta method. To keep the number of strings manageable, we introduce noise and truncate the operator $O$ at each time step by keeping the strings with the largest weight. The noise is modeled by a depolarizing channel that causes long Pauli strings to decay. In the Heisenberg picture, observables evolve under the adjoint channel. Because a depolarizing channel is self-adjoint, we can apply it directly to $O$. The transmission rate associated with a Pauli string of length $w$ is $e^{-\epsilon w}$ where $\epsilon$ is the noise amplitude. Similar approaches have been recently used in Refs.  [11, 20, 21, 22, 23, 24, 25]. The strategy here is to choose the smallest noise value that makes the simulations tractable, while not destroying the phenomena of interest.
@@ IV Krylov subspace expansion
Lets introduce Lanczos’ algorithm. The idea is to construct an orthonormal basis of operators generated by recursively applying $\mathcal{L}$ to $O$. First, one defines an inner product, which here we choose to be the Frobenius inner product (up to normalization): $AB=\frac{1}{2^{N}}{\rm Tr}[A^{\dagger}B]$ and norm $\lVert O\rVert=\frac{1}{2^{N}}{\rm Tr}[O^{2}]$. Then, the first iteration is given by:
@@ Acknowledgements.
N.L. was supported by a research grant (42085) from Villum Fonden. N.L. thanks Prof. Berislav Buca for funding and support, which made finishing this project possible. D.S. and J.C.P. are partially supported by AFOSR (Award no. FA9550-25-1-0067) and NSF (Award no. OAC-2118310). This work was supported in part through the NYU IT High Performance Computing resources, services, and staff expertise.